The displacement vectors A and B, when added together, give the resultant vector R, so that R = A + B. Use the data in the drawing and the fact that φ = 23° to find the magnitude R of the resultant vector and the angle θ that it makes with the +x axis.
1
Expert's answer
2020-12-23T10:22:17-0500




"R= \\sqrt{\n6 ^\n2\n +8 ^\n2\n +2(6)(8)cos(90+23)\u00b0\n\u200b\t\n} =7.9"

"8\cos{23°}=7.9\cosθ\\θ=\arccos0.9322=21°"